# A Survey on LLM-as-a-Judge

### v3

@@ 2.4.4. LLM-as-a-Judge for Reasoning/Thinking
The role of LLM-as-a-Judge in enhancing reasoning or thinking can be understood through two frameworks: scaling training time (Gao et al., 2023b; Trung et al., 2024) and scaling test time (Snell et al., 2024). In the training phase, LLM-as-a-Judge frequently operates within reinforcement learning paradigms, where it functions as a reward model or evaluator for data or processes. This enables the creation of high-quality reasoning datasets through mechanisms such as step-by-step verification (Lightman et al., 2023), Direct Preference Optimization(DPO) (Rafailov et al., 2024), and self-refinement (Yuan et al., 2024). Recently, several LLMs trained with reinforcement learning to exhibit advanced reasoning and thinking abilities have gained attention, such as o1222https://openai.com/index/learning-to-reason-with-llms/, DeepSeek-R1333https://api-docs.deepseek.com/news/news1120,gemini-thinking [^4] https://ai.google.dev/gemini-api/docs/thinking-mode, and QVQ [^5] https://huggingface.co/Qwen/QVQ-72B-Preview. In the test-time framework, LLM-as-a-Judge is crucial for evaluating and selecting the best reasoning paths. For example, in ”Best-of-N” generation scenarios, where multiple reasoning outputs are produced, the judge determines the most accurate and coherent response. This dual role in both training and test phases demonstrates the indispensable nature of LLM-as-a-Judge in enhancing reasoning systems.
@@ 3. Improvement Strategy
Evaluation Tasks, text width=6em [Few-shot prompting: FActScore (Min et al., 2023) / SALAD-Bench (Li et al., 2024c) / GPTScore (Fu et al., 2024)] [Evaluation steps decomposition: G-Eval (Liu et al., 2023a) / DHP (Wang et al., 2024d) / SocREval (He et al., 2024) / BSM (Saha et al., 2024)] [Evaluation criteria decomposition: HD-Eval (Liu et al., 2024a) / Hu and Gao et al. (Hu et al., 2024)] [Shuffling contents: Wang et al. (Wang et al., 2024b) / Auto-J (Li et al., 2023b) / JudgeLM (Zhu et al., 2023a) / PandaLM (Wang et al., 2023d)] [Conversion of evaluation tasks: Liu et al. (Liu et al., 2024b)] ] [Optimizing LLMs’
Output Forms, text width=6em [Constraining outputs in structured formats: G-Eval (Liu et al., 2023a) / DHP (Wang et al., 2024d) / LLM-EVAL (Lin and Chen, 2023)] [Providing evaluations with explanations: CLAIR (Chan et al., 2023) / FLEUR (Lee et al., 2024)] ] ] [Improvement Strategy
@@
LLMs’ Outputs, text width=7.5em [Score smoothing: FLEUR (Lee et al., 2024) / G-Eval (Liu et al., 2023a) / DHP (Wang et al., 2024d)] [Self validation: TrueTeacher (Gekhman et al., 2023)] ] ] ]
@@ 3.1.1. Optimizing LLMs’ Understanding of Evaluation Tasks
In addition to providing high-quality examples for LLMs, refining the evaluation task instructions is also an effective approach to optimize LLMs’ understanding of evaluation tasks. Current methods for refining evaluation tasks mainly include the decomposition of evaluation steps and criteria: (a) Decomposition of Evaluation Steps entails breaking down the entire evaluation tasks into smaller steps, providing detailed definitions and constraints for each small step in prompts, thereby guiding LLMs comprehensively through the whole evaluation pipeline. For instance, G-Eval(Liu et al., 2023a) and DHP(Wang et al., 2024d) use Chain-of-Thought(CoT)(Wei et al., 2022) to guide LLMs. SocREval(He et al., 2024) employs the Socratic method to meticulously design each step to enhance evaluation performance. Saha et al. propose Branch-Solve-Merge(BSM)(Saha et al., 2024), which divides evaluation tasks into multiple parallel sub-tasks for separate evaluation and final merge. (b) Decomposition of Evaluation Criteria involves breaking down coarse evaluation criteria like Fluency into finer-grained sub-criteria like Grammar, Engagingness, and Readability, and then generating overall scores based on these difference dimensions. HD-Eval(Liu et al., 2024a) iteratively aligns LLM evaluators with human preference via hierarchical criteria decomposition thereby addressing the potential bias in LLMs. Hu and Gao et al.(Hu et al., 2024) summarize and clearly define an explicit hierarchical classification system encompassing 11 criteria, addressing the issue of LLMs potentially confusing different evaluation standards. These refinements are specific to enable LLMs to understand the details of evaluation tasks more deeply, thereby aligning evaluation results more closely with human evaluation requirements and preferences.
Furthermore, the evaluation capabilities can be optimized based on specific shortcomings of LLMs in prompts. For instance, to address specific biases like position bias which is common in pairwise evaluations, several research efforts have optimized prompts design by randomly swapping contents to be evaluated. Wang et al.(Wang et al., 2024b) analyzed and validated the impact of position bias on LLM-as-a-judge and proposed a calibration framework to mitigate this bias by swapping the contents and averaging the scores. Auto-J(Li et al., 2023b) and JudgeLM(Zhu et al., 2023a) also enhance the evaluation consistency by shuffling the texts to be evaluated. In contrast to averaging scores, PandaLM(Wang et al., 2023d) annotates the conflicting evaluation results after swapping as ”Tie” to address the position bias.
@@ 3.1.2. Optimizing LLMs’ Output Forms
Directly requiring LLM evaluators to output evaluation results poses robustness problems. The response text may unexpectedly vary due to the inherent generative randomness of LLMs, such as outputting text like ”low relevance” while asked to measure it with discrete scores, which hinders the automated and accurate extraction of evaluation results from LLMs’ output. An effective method to enhance the robustness of output forms is to constrain LLMs’ output in structured formats within prompts. G-Eval(Liu et al., 2023a) and DHP framework(Wang et al., 2024d) perform evaluation tasks with a form-filling paradigm, constraining outputs with formats like ”X: Y”, where X represents the dimension or metric to be evaluated and Y denotes an identifiable output form like scores or specific tokens. LLM-EVAL(Lin and Chen, 2023) further modifies this form-filling paradigm, efficiently outputs evaluation results in JSON format, and obtains multidimensional scores, leveraging LLMs’ high understanding and generation capabilities of code-like textural formats.
Apart from challenges in robustness, directly outputting evaluation results by LLMs also suffer from the lack of interpretability. The meaning of evaluation results from LLM evaluators is difficult to align consistently with instructions and metrics provided in prompts. To address the challenges, CLAIR(Chan et al., 2023) requires LLMs to output evaluation scores between 0-100 simultaneously with relevant reasons as explanations in JSON format, which enhances the rationality and interpretability of the scores. FLEUR(Lee et al., 2024) utilizes LLaVA to first provide quality scores for image captions and subsequently asks with ”Why? Tell me the reason.” for explanations with the images, captions, and scores as inputs, offering a stepwise approach to provide interpretable scores.
@@ 4.2. Bias
Position Bias is the tendency of LLM evaluators to favor responses in certain positions within the prompt (Wang et al., 2024b; Thakur et al., 2024; Shi et al., 2024; Ye et al., 2024). This bias may have detrimental effects, as Vicuna-13B could outperform ChatGPT when evaluated by ChatGPT, simply by positioning the response of Vicuna-13B in the second place (Wang et al., 2024b). To measure this bias, recent work (Shi et al., 2024) proposed two metrics: Position Consistency, which quantifies how frequently a judge model selects the same response after changing their positions, and Preference Fairness, which measures the extent to which judge models favor response in certain positions. The study (Wang et al., 2024b) also introduced a metric Conflict Rate to measure the percent of disagreement after change the position of two candidate responses. Their analytical experiments reveal that the degree of positional bias fluctuates depending on the disparity in response quality and the preferred position varies with different LLMs. For instance, GPT-4 tends to favor the first position, while ChatGPT shows a preference for the second position.
@@ 4.3. Adversarial Robustness
Research (Raina et al., 2024) constructed a surrogate model from the black-box LLM-evaluator and the learn a adversarial attack phrases based on it. The evaluation score can be drastically inflated by universally inserting the learned attack phrases without improving the text quality. Furthermore, other work (Zheng et al., 2024) demonstrated that even a ”null model” that outputs a constant response irrelevant to input instructions can achieve high win rates for various LLM-as-a-judge methods. Several recent works (Ye et al., 2024; Koo et al., 2023) proposed to increase the evaluation score by adding the majority opinions, such as “90% believe this is better”. Other researches (Ye et al., 2024; Koo et al., 2023) evaluated robustness against meaningless statement in the System Prompt, e.g., “”Assistant A loves eating pasta”. These works revealed that LLM-as-a-judge are still insufficiently robust against interference irrelevant to text quality. Defensive measures like the perplexity score (Raina et al., 2024; Jain et al., 2023) can only detect limited types of adversarial examples. Therefore, constructing more robust LLM-as-a-judge is a crucial research direction for the future.
@@ 5.2.1. Comparison with Different LLMs
LLMs Alignment Biases with Position Length Concre- Empty Content Nested Familiar Human teness Reference Continuation Instruction Knowledge (n=5106) (n=2633) (n=34) (n=28) (n=26) (n=24) (n=24) (n=24) GPT-4-turbo 61.56 80.49 91.18 89.29 65.38 95.83 70.83 100.0 GPT-3.5-turbo 54.72 68.78 20.59 64.29 23.08 91.67 58.33 54.17 Qwen2.5-7B-Instruct 56.54 63.50 64.71 71.43 69.23 91.67 45.83 83.33 LLaMA3-8B-Instruct 50.72 38.85 20.59 57.14 65.38 75.00 45.83 54.17 Mistral-7B-Instruct-v0.3 55.42 59.78 26.47 67.86 53.85 66.67 37.50 41.67 Mixtral-8$\times$7B-Instruct-v0.1 56.29 59.06 50.00 78.57 42.31 83.33 29.17 83.33 gemini-2.0-thinking 60.80 76.84 94.12 89.29 50.00 100.00 83.33 100.00 o1-mini 60.19 76.73 91.18 89.29 53.85 95.83 75.00 95.83
@@
LLMs human=model1 human=model2 human=TIE (n=2070) (n=2070) (n=962) aligned accuracy (%) aligned accuracy (%) aligned accuracy (%) GPT-4-turbo 1417 68.45 1438 69.47 286 29.73 gemini-2.0-thinking 1354 65.41 1621 78.31 127 13.20 o1-mini 1444 69.76 1400 67.63 227 23.60
@@ 5.2.3. Evaluation of Reasoning LLM-as-a-Judge
As discussed in Sections 2.4 and 2.5, judgment serves as the foundation for effective reasoning capabilities. In other words, models with stronger reasoning capabilities are generally better equipped to perform as reliable judges. To validate this assumption, we conducted evaluations on several reasoning LLMs including o1-mini and Gemini-thinking. The results, presented in The results in Tables 2 and 3 yielded surprising insights: reasoning-focused LLMs failed to exhibit a clear advantage in tasks requiring alignment with human preferences. On the contrary, their performance was slightly inferior to GPT-4, suggesting that advanced reasoning does not always translate to better alignment in such tasks.

### v4

@@ 2.4.4. LLM-as-a-Judge for Reasoning/Thinking
The role of LLM-as-a-Judge in enhancing reasoning or thinking can be understood through two frameworks: scaling training time (Gao et al., 2023d; Trung et al., 2024) and scaling test time (Snell et al., 2024). In the training phase, LLM-as-a-Judge frequently operates within reinforcement learning paradigms, where it functions as a reward model or evaluator for data or processes. This enables the creation of high-quality reasoning datasets through mechanisms such as step-by-step verification (Lightman et al., 2023a), Direct Preference Optimization(DPO) (Rafailov et al., 2024), and self-refinement (Yuan et al., 2024). Recently, several LLMs trained with reinforcement learning to exhibit advanced reasoning and thinking abilities have gained attention, such as o1222https://openai.com/index/learning-to-reason-with-llms/, DeepSeek-R1333https://api-docs.deepseek.com/news/news1120,gemini-thinking [^4] https://ai.google.dev/gemini-api/docs/thinking-mode, and QVQ [^5] https://huggingface.co/Qwen/QVQ-72B-Preview. In the test-time framework, LLM-as-a-Judge is crucial for evaluating and selecting the best reasoning paths. For example, in ”Best-of-N” generation scenarios, where multiple reasoning outputs are produced, the judge determines the most accurate and coherent response. This dual role in both training and test phases demonstrates the indispensable nature of LLM-as-a-Judge in enhancing reasoning systems.
@@ 3. Improvement Strategy
Evaluation Tasks, text width=6em [Few-shot prompting: FActScore (Min et al., 2023) / SALAD-Bench (Li et al., 2024c) / GPTScore (Fu et al., 2024)] [Evaluation steps decomposition: G-Eval (Liu et al., 2023a) / DHP (Wang et al., 2024f) / SocREval (He et al., 2024) / BSM (Saha et al., 2024)] [Evaluation criteria decomposition: HD-Eval (Liu et al., 2024a) / Hu and Gao et al. (Hu et al., 2024)] [Shuffling contents: Wang et al. (Wang et al., 2024c) / Auto-J (Li et al., 2023b) / JudgeLM (Zhu et al., 2023a) / PandaLM (Wang et al., 2023d)] [Conversion of evaluation tasks: Liu et al. (Liu et al., 2024b)] ] [Optimizing LLMs’
Output Forms, text width=6em [Constraining outputs in structured formats: G-Eval (Liu et al., 2023a) / DHP (Wang et al., 2024f) / LLM-EVAL (Lin and Chen, 2023)] [Providing evaluations with explanations: CLAIR (Chan et al., 2023b) / FLEUR (Lee et al., 2024b)] ] ] [Improvement Strategy
@@
LLMs’ Outputs, text width=7.5em [Score smoothing: FLEUR (Lee et al., 2024b) / G-Eval (Liu et al., 2023a) / DHP (Wang et al., 2024f)] [Self validation: TrueTeacher (Gekhman et al., 2023)] ] ] ]
@@ 3.1.1. Optimizing LLMs’ Understanding of Evaluation Tasks
In addition to providing high-quality examples for LLMs, refining the evaluation task instructions is also an effective approach to optimize LLMs’ understanding of evaluation tasks. Current methods for refining evaluation tasks mainly include the decomposition of evaluation steps and criteria: (a) Decomposition of Evaluation Steps entails breaking down the entire evaluation tasks into smaller steps, providing detailed definitions and constraints for each small step in prompts, thereby guiding LLMs comprehensively through the whole evaluation pipeline. For instance, G-Eval(Liu et al., 2023a) and DHP(Wang et al., 2024f) use Chain-of-Thought(CoT)(Wei et al., 2022a) to guide LLMs. SocREval(He et al., 2024) employs the Socratic method to meticulously design each step to enhance evaluation performance. Saha et al. propose Branch-Solve-Merge(BSM)(Saha et al., 2024), which divides evaluation tasks into multiple parallel sub-tasks for separate evaluation and final merge. (b) Decomposition of Evaluation Criteria involves breaking down coarse evaluation criteria like Fluency into finer-grained sub-criteria like Grammar, Engagingness, and Readability, and then generating overall scores based on these difference dimensions. HD-Eval(Liu et al., 2024a) iteratively aligns LLM evaluators with human preference via hierarchical criteria decomposition thereby addressing the potential bias in LLMs. Hu and Gao et al.(Hu et al., 2024) summarize and clearly define an explicit hierarchical classification system encompassing 11 criteria, addressing the issue of LLMs potentially confusing different evaluation standards. These refinements are specific to enable LLMs to understand the details of evaluation tasks more deeply, thereby aligning evaluation results more closely with human evaluation requirements and preferences.
Furthermore, the evaluation capabilities can be optimized based on specific shortcomings of LLMs in prompts. For instance, to address specific biases like position bias which is common in pairwise evaluations, several research efforts have optimized prompts design by randomly swapping contents to be evaluated. Wang et al.(Wang et al., 2024c) analyzed and validated the impact of position bias on LLM-as-a-judge and proposed a calibration framework to mitigate this bias by swapping the contents and averaging the scores. Auto-J(Li et al., 2023b) and JudgeLM(Zhu et al., 2023a) also enhance the evaluation consistency by shuffling the texts to be evaluated. In contrast to averaging scores, PandaLM(Wang et al., 2023d) annotates the conflicting evaluation results after swapping as ”Tie” to address the position bias.
@@ 3.1.2. Optimizing LLMs’ Output Forms
Directly requiring LLM evaluators to output evaluation results poses robustness problems. The response text may unexpectedly vary due to the inherent generative randomness of LLMs, such as outputting text like ”low relevance” while asked to measure it with discrete scores, which hinders the automated and accurate extraction of evaluation results from LLMs’ output. An effective method to enhance the robustness of output forms is to constrain LLMs’ output in structured formats within prompts. G-Eval(Liu et al., 2023a) and DHP framework(Wang et al., 2024f) perform evaluation tasks with a form-filling paradigm, constraining outputs with formats like ”X: Y”, where X represents the dimension or metric to be evaluated and Y denotes an identifiable output form like scores or specific tokens. LLM-EVAL(Lin and Chen, 2023) further modifies this form-filling paradigm, efficiently outputs evaluation results in JSON format, and obtains multidimensional scores, leveraging LLMs’ high understanding and generation capabilities of code-like textural formats.
Apart from challenges in robustness, directly outputting evaluation results by LLMs also suffer from the lack of interpretability. The meaning of evaluation results from LLM evaluators is difficult to align consistently with instructions and metrics provided in prompts. To address the challenges, CLAIR(Chan et al., 2023b) requires LLMs to output evaluation scores between 0-100 simultaneously with relevant reasons as explanations in JSON format, which enhances the rationality and interpretability of the scores. FLEUR(Lee et al., 2024b) utilizes LLaVA to first provide quality scores for image captions and subsequently asks with ”Why? Tell me the reason.” for explanations with the images, captions, and scores as inputs, offering a stepwise approach to provide interpretable scores.
@@ 4.2. Bias
Position Bias is the tendency of LLM evaluators to favor responses in certain positions within the prompt (Wang et al., 2024c; Thakur et al., 2024; Shi et al., 2024; Ye et al., 2024). This bias may have detrimental effects, as Vicuna-13B could outperform ChatGPT when evaluated by ChatGPT, simply by positioning the response of Vicuna-13B in the second place (Wang et al., 2024c). To measure this bias, recent work (Shi et al., 2024) proposed two metrics: Position Consistency, which quantifies how frequently a judge model selects the same response after changing their positions, and Preference Fairness, which measures the extent to which judge models favor response in certain positions. The study (Wang et al., 2024c) also introduced a metric Conflict Rate to measure the percent of disagreement after change the position of two candidate responses. Their analytical experiments reveal that the degree of positional bias fluctuates depending on the disparity in response quality and the preferred position varies with different LLMs. For instance, GPT-4 tends to favor the first position, while ChatGPT shows a preference for the second position.
@@ 4.3. Adversarial Robustness
Research (Raina et al., 2024) constructed a surrogate model from the black-box LLM-evaluator and the learn a adversarial attack phrases based on it. The evaluation score can be drastically inflated by universally inserting the learned attack phrases without improving the text quality. Similarly, work by Lee et al. (Lee et al., 2024a) introduced EMBER, a benchmark that revealed biases in when assessing outputs with epistemic markers, such as expressions of certainty or uncertainty. Furthermore, other work (Zheng et al., 2024a) demonstrated that even a ”null model” that outputs a constant response irrelevant to input instructions can achieve high win rates for various LLM-as-a-judge methods. Several recent works (Ye et al., 2024; Koo et al., 2023) proposed to increase the evaluation score by adding the majority opinions, such as “90% believe this is better”. Other research (Ye et al., 2024; Koo et al., 2023) evaluated robustness against meaningless statement in the System Prompt, e.g., “”Assistant A loves eating pasta”. These works revealed that LLM-as-a-judge are still insufficiently robust against interference irrelevant to text quality. Defensive measures like the perplexity score (Raina et al., 2024; Jain et al., 2023) can only detect limited types of adversarial examples. Therefore, constructing more robust LLM-as-a-judge is a crucial research direction for the future.
@@ 5.2.1. Comparison with Different LLMs
LLMs Alignment Biases with Position Length Concre- Empty Content Nested Familiar Human teness Reference Continuation Instruction Knowledge (n=5106) (n=2633) (n=34) (n=28) (n=26) (n=24) (n=24) (n=24) GPT-4-turbo 61.54 80.31 91.18 89.29 65.38 95.83 70.83 100.0 GPT-3.5-turbo 54.72 68.78 20.59 64.29 23.08 91.67 58.33 54.17 Qwen2.5-7B-Instruct 56.54 63.50 64.71 71.43 69.23 91.67 45.83 83.33 LLaMA3-8B-Instruct 50.72 38.85 20.59 57.14 65.38 75.00 45.83 54.17 Mistral-7B-Instruct-v0.3 55.42 59.78 26.47 67.86 53.85 66.67 37.50 41.67 Mixtral-8$\times$7B-Instruct-v0.1 56.29 59.06 50.00 78.57 42.31 83.33 29.17 83.33 gemini-2.0-thinking 60.75 76.84 94.12 89.29 50.00 100.00 83.33 100.00 o1-mini 60.16 76.73 91.18 89.29 53.85 95.83 75.00 95.83 deepseek r1 56.48 69.17 94.12 100.00 50.00 100.00 75.00 87.50
@@
LLMs human=model1 human=model2 human=TIE (n=2071) (n=2071) (n=964) aligned accuracy (%) aligned accuracy (%) aligned accuracy (%) GPT-4-turbo 1418 68.47 1438 69.44 286 29.67 gemini-2.0-thinking 1354 65.38 1621 78.27 127 13.17 o1-mini 1444 69.72 1401 67.65 227 23.55 deepseek r1 1369 66.10 1342 64.80 173 17.95
@@ 5.2.3. Evaluation of Reasoning LLM-as-a-Judge
As discussed in Sections 2.4 and 2.5, judgment serves as the foundation for effective reasoning capabilities. In other words, models with stronger reasoning capabilities are generally better equipped to perform as reliable judges. To validate this assumption, we conducted evaluations on several reasoning LLMs including o1-mini, Gemini-thinking and Deepseek-R1. The results in Tables 2 and 3 provide key insights into the performance of reasoning-focused LLMs. While these models—gemini-2.0-thinking, o1-mini, and deepseek r1—demonstrate competitive alignment and accuracy relative to the top-performing GPT-4-turbo, their improvements in tasks requiring human alignment are not as pronounced as expected.
GPT-4-turbo remains the benchmark for alignment, achieving the highest accuracy rates of 68.47% and 69.44% in the human=model1 and human=model2 scenarios, respectively. It also excels in resolving ambiguous cases, with an accuracy of 29.67% in the human=TIE scenario, outperforming all other models. Among reasoning-enhanced models, gemini-2.0-thinking shows strong performance in the human=model2 scenario, achieving an accuracy of 78.27%—surpassing GPT-4-turbo—highlighting its capability to identify models that align more closely with human judgment. However, this strength does not consistently extend to other tasks or scenarios. Similarly, o1-mini and deepseek-R1, while trailing slightly behind GPT-4-turbo and gemini-2.0-thinking, outperform non-GPT models like Mixtral-8$\times$7B-Instruct-v0.1, demonstrating the added value of reasoning-based enhancements in alignment tasks. These results indicate that reasoning-enhanced LLMs provide meaningful advancements over baseline models but fall short of delivering consistent advantages in alignment-related tasks, suggesting room for further optimization in this area.
@@ 5.2.4. Summary
6. LLM-as-a-Judge and o1-like Reasoning Enhancement
When faced with a challenging problem, humans often spend considerable time and effort reflecting on various possibilities before arriving at a solution. In a similar fashion, o1, an advanced model developed by OpenAI, engages in a structured chain of thought to solve complex tasks. This process of deliberate reasoning allows o1 to continuously refine its approach (Qin et al., 2024c), step by step, as it navigates through difficult scenarios. A key factor in enhancing o1’s reasoning is the integration of LLM-as-a-Judge, which evaluates the model’s reasoning paths at each stage. As o1 works through a problem, the judge provides feedback that helps the model improve by pointing out inconsistencies, suggesting corrections, and identifying simpler ways to break down difficult tasks. By using the feedback from its own evaluations, much like a constitutional AI framework, o1 is able to adapt its reasoning strategies and enhance its performance. Through reinforcement learning, o1 fine-tunes its strategies, learning not only from its successes but also from its mistakes. The combination of LLM-as-a-Judge, reinforcement learning, and the feedback loop from constitutional evaluation allows o1 to dynamically adjust its reasoning, ensuring that the model continuously improves its ability to solve complex problems over time. This synergy between reasoning and judgment, combined with continuous feedback, drives o1’s advanced problem-solving capabilities.
Figure 14. The relationship of LLM-as-a-Judge and Reasoning/Thinking.
In this process, we can observe two ways in which LLM-as-a-Judge evaluates both Reasoning and Thinking. The first method involves evaluating the reasoning process during the training phase, where LLM-as-a-Judge provides feedback that is used to fine-tune the model through reinforcement learning, enhancing its reasoning ability. This feedback helps o1 refine its approach, identify errors, and break down complex tasks into more manageable components. The second method occurs during test time, where LLM-as-a-Judge dynamically evaluates the model’s reasoning outputs, providing real-time feedback that further improves the model’s performance. What both methods share is their ability to offer o1 continuous feedback, whether it is positive or negative, which drives a process of self-improvement. By incorporating this feedback into its reasoning process, o1 is able to iteratively adjust its approach and learn from its mistakes. This cycle of reflection and correction enhances the model’s ability to think critically and solve increasingly complex problems. The synergy between the two evaluation strategies—during both training and testing—creates a powerful feedback loop that allows o1 to dynamically optimize its reasoning and thinking capabilities over time, leading to significant advancements in its problem-solving abilities.
Constitutional AI (Bai et al., 2022), which was used to build DeepSeek-R1 can be seen as a specific form of LLM-as-a-Judge, where the model uses its own evaluations, such as voting results, as feedback to guide its optimization. In this approach, o1 evaluates its reasoning through internal assessments, refining its decisions based on predefined principles. This self-generated feedback loop helps o1 correct errors and improve its performance over time, without needing external verification. By integrating LLM-as-a-Judge within the Constitutional AI framework, o1 continuously adjusts its reasoning strategies, leading to better problem-solving abilities through self-improvement and reinforcement learning.
Relationship Between LLM-as-a-Judge and Reasoning
Reasoning, as a cognitive process, involves applying logic and evidence to derive conclusions. It is central to intellectual tasks such as decision-making, problem-solving, and critical analysis. Reasoning requires evaluating multiple possibilities and determining the most logically sound and coherent path. In contrast, LLM-as-a-Judge refers to the use of LLMs to perform judgment tasks, such as evaluating, scoring, ranking, or selecting optimal answers based on generated outputs. This concept parallels a judge’s role in ensuring fairness, accuracy, and coherence in competitive contexts. Although reasoning and judgment are separate concepts, they are closely connected. Reasoning frequently relies on judgment to evaluate intermediate steps, improve logic, and guarantee clarity in its results. When the process involves an infinite number of judgments, we can consider it as a process that approximates both reasoning and thinking. At the same time, effective judgment relies on strong reasoning capabilities to evaluate options against a set of logical criteria, which we discussed in Section 2.5 and Section 5.2.3. Therefore, LLM-as-a-Judge not only evaluates outputs but also enhances the reasoning process by helping to identify the most coherent, accurate solutions.